OpenStudy (anonymous):

An equation is shown below: 5(2x - 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

OpenStudy (playgames1122):

10x-15=5 10x = 5+15 x= 20/10 x=2

OpenStudy (playgames1122):

So its part A

OpenStudy (anonymous):

Blahahahahahahahah